OpenStudy (beststudent):

A person expends $240 in the purchase of wheat. If he had paid 20 cents a bushel less he could have obtained 100 bushels more for the same money. How many bushels did he buy?

OpenStudy (mathstudent55):

He bought x bushels of wheat. The price per bushel was y. xy = 240 Ok so far?

OpenStudy (mathstudent55):

Now use again x as the number of bushels he actually bought and y as the actual cost per bushel and write an equation for "If he had paid 20 cents a bushel less he could have obtained 100 bushels more for the same money."

OpenStudy (mathstudent55):

Do you follow?

OpenStudy (beststudent):

Yes. Continue please?

OpenStudy (beststudent):

But I'm not sure how to form an equation for "If he had paid 20 cents a bushel less he could have obtained 100 bushels more for the same money."

OpenStudy (mathstudent55):

Ok. Let's work on that together. First, keep this in mind: The number of bushels multiplied by the cost per bushel = the total price paid. For example, if a bushel costs $1.25, and he bought 100 bushels, then the total cost is 100 bushels * 1.25 dollars/bushel = $125

OpenStudy (beststudent):

Yes, I got that.

OpenStudy (mathstudent55):

Above we used x as the number of bushels. y is the cost per bushel. That is why xy (x multiplied by y) is the total price he paid. Now let's look at what he could have done. The price could have been $0.20 less per bushel. If the price per bushel is y, how do you write 20 cents less than y as an expression using y?

OpenStudy (beststudent):

y-0.2 ?

OpenStudy (mathstudent55):

Correct. If the price had been $0.20 less per bushel, he would have bought 100 more bushels. He bought x bushels. How do you write 100 more than x as an expression using x.

OpenStudy (beststudent):

x+100

OpenStudy (mathstudent55):

Correct again.

OpenStudy (mathstudent55):

Now you have a new cost per bushel: y = 0.2, and a new number of bushels: x + 100 The total cost remains $240. Can you write an expression that shows the new cost per bushel times the new number of bushels equaling the total cost?

OpenStudy (beststudent):

(y-0.2)(x+100)=240

OpenStudy (beststudent):

Is the answer 300 bushels?

OpenStudy (mathstudent55):

Correct. That equation, together with xy = 240, forms a system of equations. Now you need to solve the system of equations to find x and y.

OpenStudy (beststudent):

Does x=300 and x= -400? Therefore, the answer is 300 because it has to be positive?

OpenStudy (mathstudent55):

Let's see: xy = 240 Solve for y: y = 240/x Substitute in other equation: (240/x - 0.2)(x +100) = 240 240 + 24000/x - 0.2x - 20 = 240 24000/x - 0.2x - 20 = 0 24000 - 0.2x^2 - 20x = 0 0.2x^2 + 20x - 24000 = 0 x^2 + 100x - 120,000 = 0 (x + 400)(x - 300) = 0 x + 400 = 0 or x - 300 = 0 x = 400 or x = -300 You are correct.

OpenStudy (beststudent):

Isn't -400 and 300?

OpenStudy (mathstudent55):

We can check. Since x = 300, then xy = 240, giving us y = 240/x = 240/300 = 0.8 The original cost was $0.80 per bushel.

OpenStudy (mathstudent55):

Yes, of course, I just copied it wrong.

OpenStudy (mathstudent55):

Look at my next line where I used x = 300.

OpenStudy (mathstudent55):

The original cost was $0.80 per bushel. Then $0.80 * 300 = $240 Now do $0.20 less per bushel, and 100 bushels more: $0.60 * 400 = $240 Therefore, our answer of 300 bushels is correct.

OpenStudy (mathstudent55):

Great work!

OpenStudy (beststudent):

Thank you!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (beststudent):

I'm not sure but my teacher said the answer is 120,000. Is there a possible way for that to happen? My teacher is going to go over it next week, but I hate waiting!

OpenStudy (beststudent):

@mathstudent55

OpenStudy (mathstudent55):

The answer to this problem is not 120,000. If $240 bought 120,000 bushels, then each bushel cost $0.002, or 0.2 cents. It would then be impossible to pay $0.20 less per bushel. Now look at the answer we got. The number of bushels bought was 300. Each bushel cost $240/300 = $0.80 Now reduce the price of each bushel by $0.20 to have a per bushel price of $0.60. The same $240 now buys you 240/0.60 = 400 bushels, which is 100 more bushels than 300, just as the problem stated.